OpenStudy (anonymous):

what is the pattern for the sequence 3 ,12, 27, 48?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

75 is the next

OpenStudy (anonymous):

t(n) = 3n^2

OpenStudy (anonymous):

how did u get it

OpenStudy (anonymous):

the adding goes 9, 15, 21 (12-3=9. 27-12=15) but in between each is +6 (ie. 15-9 = 6. 9+6=15, etc.) so the next would be 54.

OpenStudy (anonymous):

9,15,21, etc.

OpenStudy (anonymous):

how did you get 3n^2?

OpenStudy (anonymous):

by observation?

OpenStudy (anonymous):

can you explain?

OpenStudy (anonymous):

how can u explain observation?